Which of the following simplifies to a negative number?
A.)-4^-4
B.)(-4)^-4
C.)4^-4
D.)1/4^-4
please help.
Thank you.

Answer by stanbon(75887)

Which of the following simplifies to a negative number?
A.)-4^-4 = -(4)^-4 = -1/256
------------------------------
B.)(-4)^-4 = (1/(-4)^4) = 1/256
------------------------------
C.)4^-4 = 1/256
-----------------------
D.)(1/4)^-4 = 4^4 = 256
===========================
Cheers,
Stan H.
==========================